# System-2 Reasoning via Generality and Adaptation

### v1

@@ Abstract
While significant progress has been made in task-specific applications, current models struggle with deep reasoning, generality, and adaptation—key components of System-2 reasoning that are crucial for achieving Artificial General Intelligence (AGI). Despite the promise of approaches such as program synthesis, language models, and transformers, these methods often fail to generalize beyond their training data and to adapt to novel tasks, limiting their ability to perform human-like reasoning. This paper explores the limitations of existing approaches in achieving advanced System-2 reasoning and highlights the importance of generality and adaptation for AGI. Moreover, we propose four key research directions to address these gaps: (1) learning human intentions from action sequences, (2) combining symbolic and neural models, (3) meta-learning for unfamiliar environments, and (4) reinforcement learning to reason multi-step. Through these directions, we aim to advance the ability to generalize and adapt, bringing computational models closer to the reasoning capabilities required for AGI.
@@ 1 Introduction
Nowadays, AI has made significant strides in task-specific applications, mainly through the success of neural networks [15, 28, 41, 61]. Furthermore, AI models based on language models [1, 6] and reinforcement learning models [25, 51] are achieving human-level performance with surprising results on several tasks. However, achieving human-level deep and logical reasoning, referred to as System-2 reasoning [29], a concept central to human cognition and intelligence, remains an open challenge. System-2 reasoning is characterized by abstract thought, logical deduction, and the ability to adapt to novel and complex situations, critical components of AGI.
@@
There is growing recognition that for AI to meet the reasoning requirements of AGI, it must excel in generality and adaptation [24], which are two fundamental capabilities that enable models to handle unseen tasks and unpredictable environments. Generality helps AI apply learned knowledge to new contexts, while adaptation enables flexible responses to changing scenarios [14]. These capabilities are also crucial for System-2 reasoning because they are fundamental to human-like intelligence [33].
@@
Despite recent advancements, current AI models are limited by their reliance on training data and task-specific optimization [57]. Neural networks such as Large Language Models (LLMs) [17, 23, 36], and Model-Based RL (MBRL) models [51], while highly capable in constrained settings, often fail to generalize effectively beyond their training distributions. This has led to a critical gap between existing AI models and the level of reasoning required for AGI. In particular, most models demonstrate System-1 reasoning, which refers to fast, pattern-based decision-making, while lacking the deliberative and adaptable qualities inherent in System-2 reasoning [5, 11].
@@
To bridge this gap, it is essential to identify strategies that foster both generality and adaptation in AI models. The Abstraction and Reasoning Corpus (ARC) [14] has been proposed as a benchmark specifically designed to assess an AI model’s ability to generalize and reason abstractly in unfamiliar scenarios, without relying on extensive training data. Previous approaches to solving ARC, such as program synthesis [2, 3, 4, 8, 9, 20, 26, 37, 38, 50, 65], LLMs [13, 22, 27, 40, 48, 53, 59, 62, 63, 66], and variation of transformers [7, 44, 46] have shown promise but continue to struggle with abstract reasoning and adaptability when applied to novel tasks. For AI to evolve toward AGI-level reasoning, we must advance beyond current methodologies and explore new directions that focus on enhancing System-2 reasoning.
@@
This paper addresses the limitations of existing AI methods in achieving System-2 reasoning and proposes four key research directions: (1) learning human intentions from trajectories, (2) integrating symbolic and neural hybrid models, (3) employing meta-learning for adaptation, and (4) using reinforcement learning to enhance multi-step reasoning. In this paper, we explore each key research direction for improving AI’s System-2 reasoning capabilities and demonstrate how these abilities can be assessed using benchmarks such as the Abstraction and Reasoning Corpus (ARC) [14], which focuses on generality and adaptation. By exploring these directions, we aim to improve AI’s generality and adaptability, bringing AI models closer to the reasoning capabilities required for AGI.
@@ 2 Backgrounds
2.1 Two Modes of Reasoning: System-1 and System-2
The concepts of System-1 and System-2 reasoning, introduced in psychology, describe two distinct modes of thinking [29]. System-1 is fast and intuitive, relying on automatic decision-making based on pattern recognition. Humans use System-1 in everyday situations that require little cognitive effort, such as driving or performing simple arithmetic, where decisions are made quickly based on learned experiences. In AI, this can be emulated using neural networks that learn rapidly and, when combined with tree search, replicate fast decision-making [5]. However, System-1 models work well in familiar environments but struggle with new situations and complex tasks.
On the other hand, System-2 reasoning involves slower, more deliberate thought processes that require logical deduction and abstract thinking. In contrast to System-1, System-2 is needed when encountering novel tasks that demand reasoning and planning, rather than relying on pattern recognition. This type of reasoning is essential for solving new challenges and adapting to dynamic environments, as is often the case when humans develop strategies for unfamiliar tasks. Introducing System-2 reasoning into AI would address current AI models’ limitations in handling more abstract, logic-driven tasks and enhance their ability to reason more like humans [11].
System-1 and System-2 differ in their cognitive mechanisms. System-1 is fast, intuitive, and pattern-based, while System-2 is slow, deliberate, and relies on logical reasoning. System-1 operates efficiently in familiar situations but lacks the flexibility to adapt to new environments, whereas System-2 is crucial for handling more complex, unfamiliar tasks. In AGI research, the integration of both systems is essential for achieving human-like reasoning capabilities, enabling AI to not only replicate learned behaviors but also think abstractly and adapt to novel challenges.
@@ 2.2 Generality and Adaptation: Two Key Components of AGI Reasoning
While System-2 reasoning enables deep, logical thought processes for handling complex and novel tasks, two critical components are required to fully realize its potential in AI models: generality and adaptation. For AI to achieve human-like intelligence, it must not only excel in task-specific skills but also demonstrate the ability to generalize and adapt across a wide range of environments and tasks. These two capabilities, generality and adaptation, are fundamental components of Artificial General Intelligence (AGI). Generality enables AI models to apply learned knowledge to novel contexts, allowing them to operate in unfamiliar situations without extensive retraining. Adaptation refers to the ability of AI to modify its behavior in response to changing environments, ensuring that it can handle new tasks and challenges as they arise.
@@ 2.3 ARC: A Benchmark for Measuring Generality and Adaptation
ARC is specifically designed to assess two core components of AGI, generality and adaptation. For an AI model to succeed in ARC tasks, it must generalize from limited input-output examples and adapt its strategies to solve new, unseen tasks. Thus, ARC challenges AI models to demonstrate both the ability to apply learned knowledge in novel contexts (generality) and the capacity to adjust to varying conditions or rules (adaptation). ARC is not only a benchmark for generality and adaptation but also serves as a comprehensive test for System-2 reasoning in AI. The tasks in ARC require the type of generality and adaptation that are central to System-2 reasoning, making it an ideal tool for evaluating the research directions proposed in this paper. By succeeding in ARC tasks, AI models can demonstrate progress towards achieving System-2 reasoning, which is a crucial step towards AGI.
@@ 3 Analysis of Existing Approaches
In addressing ARC tasks, several approaches in AI have been explored, including program synthesis, LLMs, and transformers. In the view of System-2 reasoning, while these approaches have advanced generalization, they still face difficulties when adapting to novel and unfamiliar tasks. By reviewing these methods, we aim to highlight their successes and identify areas where further research is needed to improve both generality and adaptation.
@@ 3.3 Transformers Also Enhance Generality but Struggle with Adaptation
4 Proposed Research Directions Toward System-2 Reasoning
In the previous section, we explored how generality and adaptation are critical components of System-2 reasoning. While existing AI models have shown mixed results in terms of generality, they consistently struggle with adaptation, which limits their ability to handle novel and complex tasks. To address these limitations, this section proposes four research directions aimed at advancing System-2 reasoning. AI must be able to extract abstract knowledge (), such as human intentions, from data and combine this knowledge in a generalized way, as seen in neuro-symbolic approaches. Additionally, meta-learning plays a crucial role in enabling AI to adapt efficiently across diverse environments and tasks, building on the foundation of learned models. Finally, reinforcement learning serves as a framework that integrates these research directions, providing the necessary structure to improve both generality and adaptation.
@@ 4.1 Learning Human Intentions from Data Supports Logical Deduction
To achieve System-2 reasoning, it is essential to establish a foundation for logical deduction and abstract thinking by capturing human intentions. Humans inherently perform System-2 reasoning through logical and abstract processes, and human intentions contain these crucial elements. Therefore, learning from human intentions, rather than merely imitating human actions, can provide a strong basis for advancing System-2 reasoning. Benchmarks like ARC [14], mini-ARC [30], 1D-ARC [66], ConceptARC [42], LARC [2], and MC-ARC [53] enable the collection of human behavior data to train AI models. Platforms like ARCreate Playground [31], O2ARC [52], ARC-Interactive [56], and ARC-Game [12] are useful for collecting such data. Additionally, environments like ARCLE [34] provide settings where models can learn from the collected data to improve reasoning and adaptability.
@@
These approaches are particularly effective for learning intentions from human trajectories in ARC tasks, where understanding and adapting to human strategies are crucial for success. For instance, in a task where the objective is to align a grid diagonally, Topic Modeling could help AI identify repeated rotations and flips as strong indicators of diagonal alignment strategies used by humans. Sequential Pattern Mining can further enhance this understanding by detecting recurring action sequences, such as Rotate(CW) followed by Flip(Horizontal), that consistently lead to correct outcomes. By identifying these patterns, the AI can begin to predict successful strategies and apply them to other, similarly structured tasks. Hidden Markov Models take this reasoning process a step further by modeling the transitions between actions and hidden goals, allowing the AI to generalize learned strategies across a wider variety of novel tasks. Incorporating these methods enables AI models to develop a deeper understanding of human reasoning, enhancing their logical deduction and adaptability, which are fundamental components of achieving System-2 reasoning capabilities. Thus, the combination of these techniques with data collected through platforms like O2ARC [52] not only improves AI’s ability to perform abstract reasoning but also strengthens its potential for handling complex, multi-step tasks in dynamic environments.
@@ 4.2 Combining Symbolic and Neural Models Improves Abstraction and Generality
Achieving System-2 reasoning requires models to move beyond simple pattern recognition and incorporate logical, structured reasoning capabilities. One promising approach to achieving this is through hybrid models that combine neural networks with symbolic reasoning. Neural networks excel at learning patterns from data, while symbolic reasoning provides interpretable, rule-based frameworks for logical deduction. By integrating these two methodologies, it is possible to enhance the ability to reason abstractly and generalize to new, unseen tasks, which are essential for System-2 reasoning.
Several methods have been developed to combine symbolic and neural models, enhancing the generality and reasoning abilities.
Neuro-Symbolic Programming (NSP) [45] integrates the pattern recognition strengths of neural networks with the explicit logic of symbolic programming. In this approach, the neural network learns patterns from data, while the symbolic component generates interpretable code that can be applied to reasoning tasks. This hybrid method enables handling of complex tasks requiring both learning from examples and reasoning over abstract concepts, such as program synthesis and abstract reasoning.
@@
In ARC tasks, hybrid neuro-symbolic models are well-suited for generalizing from a small number of examples and applying symbolic rules to solve new tasks. For example, NSP can learn how humans apply transformations like rotations and flips to grids, and then generalize these transformations using symbolic logic. ILP could infer rules based on observed task patterns, while symbolic regression might identify underlying mathematical relationships between grid elements. By combining neural networks and symbolic reasoning, these hybrid models can reason more effectively about the structure of tasks, enhancing their abstract reasoning capabilities and enabling more generalize to various tasks, key aspects of System-2 reasoning.
@@ 4.3 Meta-Learning Facilitates Adaptation to Unfamiliar Environments
System-2 reasoning requires AI systems to adapt quickly to new and unfamiliar tasks, a process that can be supported by meta-learning. Unlike traditional learning methods that are focused on task-specific optimization, meta-learning trains models to "learn how to learn,"making them more adaptable to various environments. By facilitating faster adaptation and improved generalization, meta-learning methods enable AI systems to respond effectively to novel challenges, which is crucial for achieving System-2 reasoning in tasks that require flexible thinking and logical deduction.
@@
Prototypical Networks [54] are designed for few-shot learning, where models classify new examples by comparing them to prototype representations. By learning these prototypes, AI systems can generalize across tasks with minimal data, making them efficient in environments that require structural reasoning. Prototypical networks are especially beneficial for System-2 reasoning, where recognizing and adapting to new patterns is key.
@@
In the context of ARC tasks, where multiple transformations may be required to achieve a goal, meta-learning can significantly enhance adaptability. For instance, in a task where the objective is to align objects diagonally, RNP could track the sequence of transformations and adjust reasoning based on temporal patterns. Prototypical networks could help the system recognize structural patterns involved in diagonal alignment, facilitating generalization across tasks. Meta-RL would provide the framework for balancing long-term planning with real-time reasoning, enabling the system to dynamically adjust its strategy based on feedback. By leveraging meta-learning, AI models can enhance their reasoning and adaptation capabilities, bringing them closer to the level of System-2 reasoning.

### v3

# System 2 Reasoning via Generality and Adaptation

@@ Abstract
While significant progress has been made in task-specific applications, current models struggle with deep reasoning, generality, and adaptation—key components of System 2 reasoning that are crucial for achieving Artificial General Intelligence (AGI). Despite the promise of approaches such as program synthesis, language models, and transformers, these methods often fail to generalize beyond their training data and to adapt to novel tasks, limiting their ability to perform human-like reasoning. This paper explores the limitations of existing approaches in achieving advanced System 2 reasoning and highlights the importance of generality and adaptation for AGI. Moreover, we propose four key research directions to address these gaps: (1) learning human intentions from action sequences, (2) combining symbolic and neural models, (3) meta-learning for unfamiliar environments, and (4) reinforcement learning to reason multi-step. Through these directions, we aim to advance the ability to generalize and adapt, bringing computational models closer to the reasoning capabilities required for AGI.
@@ 1 Introduction
Nowadays, AI has made significant strides in task-specific applications, mainly through the success of neural networks [15, 28, 41, 61]. Furthermore, AI models based on language models [1, 6] and reinforcement learning models [25, 51] are achieving human-level performance with surprising results on several tasks. However, achieving human-level deep and logical reasoning, referred to as System 2 reasoning [29], a concept central to human cognition and intelligence, remains an open challenge. System 2 reasoning is characterized by abstract thought, logical deduction, and the ability to adapt to novel and complex situations, critical components of AGI.
@@
There is growing recognition that for AI to meet the reasoning requirements of AGI, it must excel in generality and adaptation [24], which are two fundamental capabilities that enable models to handle unseen tasks and unpredictable environments. Generality helps AI apply learned knowledge to new contexts, while adaptation enables flexible responses to changing scenarios [14]. These capabilities are also crucial for System 2 reasoning because they are fundamental to human-like intelligence [33].
@@
Despite recent advancements, current AI models are limited by their reliance on training data and task-specific optimization [57]. Neural networks such as Large Language Models (LLMs) [17, 23, 36], and Model-Based RL (MBRL) models [51], while highly capable in constrained settings, often fail to generalize effectively beyond their training distributions. This has led to a critical gap between existing AI models and the level of reasoning required for AGI. In particular, most models demonstrate System 1 reasoning, which refers to fast, pattern-based decision-making, while lacking the deliberative and adaptable qualities inherent in System 2 reasoning [5, 11].
@@
To bridge this gap, it is essential to identify strategies that foster both generality and adaptation in AI models. The Abstraction and Reasoning Corpus (ARC) [14] has been proposed as a benchmark specifically designed to assess an AI model’s ability to generalize and reason abstractly in unfamiliar scenarios, without relying on extensive training data. Previous approaches to solving ARC, such as program synthesis [2, 3, 4, 8, 9, 20, 26, 37, 38, 50, 65], LLMs [13, 22, 27, 40, 48, 53, 59, 62, 63, 66], and variation of transformers [7, 44, 46] have shown promise but continue to struggle with abstract reasoning and adaptability when applied to novel tasks. For AI to evolve toward AGI-level reasoning, we must advance beyond current methodologies and explore new directions that focus on enhancing System 2 reasoning.
@@
This paper addresses the limitations of existing AI methods in achieving System 2 reasoning and proposes four key research directions: (1) learning human intentions from trajectories, (2) integrating symbolic and neural hybrid models, (3) employing meta-learning for adaptation, and (4) using reinforcement learning to enhance multi-step reasoning. In this paper, we explore each key research direction for improving AI’s System 2 reasoning capabilities and demonstrate how these abilities can be assessed using benchmarks such as the Abstraction and Reasoning Corpus (ARC) [14], which focuses on generality and adaptation. By exploring these directions, we aim to improve AI’s generality and adaptability, bringing AI models closer to the reasoning capabilities required for AGI.
@@ 2 Backgrounds
2.1 Two Modes of Reasoning: System 1 and System 2
The concepts of System 1 and System 2 reasoning, introduced in psychology, describe two distinct modes of thinking [29]. System 1 is fast and intuitive, relying on automatic decision-making based on pattern recognition. Humans use System 1 in everyday situations that require little cognitive effort, such as driving or performing simple arithmetic, where decisions are made quickly based on learned experiences. In AI, this can be emulated using neural networks that learn rapidly and, when combined with tree search, replicate fast decision-making [5]. However, System 1 models work well in familiar environments but struggle with new situations and complex tasks.
On the other hand, System 2 reasoning involves slower, more deliberate thought processes that require logical deduction and abstract thinking. In contrast to System 1, System 2 is needed when encountering novel tasks that demand reasoning and planning, rather than relying on pattern recognition. This type of reasoning is essential for solving new challenges and adapting to dynamic environments, as is often the case when humans develop strategies for unfamiliar tasks. Introducing System 2 reasoning into AI would address current AI models’ limitations in handling more abstract, logic-driven tasks and enhance their ability to reason more like humans [11].
System 1 and System 2 differ in their cognitive mechanisms. System 1 is fast, intuitive, and pattern-based, while System 2 is slow, deliberate, and relies on logical reasoning. System 1 operates efficiently in familiar situations but lacks the flexibility to adapt to new environments, whereas System 2 is crucial for handling more complex, unfamiliar tasks. In AGI research, the integration of both systems is essential for achieving human-like reasoning capabilities, enabling AI to not only replicate learned behaviors but also think abstractly and adapt to novel challenges.
@@ 2.2 Generality and Adaptation: Two Key Components of AGI Reasoning
While System 2 reasoning enables deep, logical thought processes for handling complex and novel tasks, two critical components are required to fully realize its potential in AI models: generality and adaptation. For AI to achieve human-like intelligence, it must not only excel in task-specific skills but also demonstrate the ability to generalize and adapt across a wide range of environments and tasks. These two capabilities, generality and adaptation, are fundamental components of Artificial General Intelligence (AGI). Generality enables AI models to apply learned knowledge to novel contexts, allowing them to operate in unfamiliar situations without extensive retraining. Adaptation refers to the ability of AI to modify its behavior in response to changing environments, ensuring that it can handle new tasks and challenges as they arise.
@@ 2.3 ARC: A Benchmark for Measuring Generality and Adaptation
ARC is specifically designed to assess two core components of AGI, generality and adaptation. For an AI model to succeed in ARC tasks, it must generalize from limited input-output examples and adapt its strategies to solve new, unseen tasks. Thus, ARC challenges AI models to demonstrate both the ability to apply learned knowledge in novel contexts (generality) and the capacity to adjust to varying conditions or rules (adaptation). ARC is not only a benchmark for generality and adaptation but also serves as a comprehensive test for System 2 reasoning in AI. The tasks in ARC require the type of generality and adaptation that are central to System 2 reasoning, making it an ideal tool for evaluating the research directions proposed in this paper. By succeeding in ARC tasks, AI models can demonstrate progress towards achieving System 2 reasoning, which is a crucial step towards AGI.
@@ 3 Analysis of Existing Approaches
In addressing ARC tasks, several approaches in AI have been explored, including program synthesis, LLMs, and transformers. In the view of System 2 reasoning, while these approaches have advanced generalization, they still face difficulties when adapting to novel and unfamiliar tasks. By reviewing these methods, we aim to highlight their successes and identify areas where further research is needed to improve both generality and adaptation.
@@ 3.3 Transformers Also Enhance Generality but Struggle with Adaptation
4 Proposed Research Directions Toward System 2 Reasoning
In the previous section, we explored how generality and adaptation are critical components of System 2 reasoning. While existing AI models have shown mixed results in terms of generality, they consistently struggle with adaptation, which limits their ability to handle novel and complex tasks. To address these limitations, this section proposes four research directions aimed at advancing System 2 reasoning. AI must be able to extract abstract knowledge (), such as human intentions, from data and combine this knowledge in a generalized way, as seen in neuro-symbolic approaches. Additionally, meta-learning plays a crucial role in enabling AI to adapt efficiently across diverse environments and tasks, building on the foundation of learned models. Finally, reinforcement learning serves as a framework that integrates these research directions, providing the necessary structure to improve both generality and adaptation.
@@ 4.1 Learning Human Intentions from Data Supports Logical Deduction
To achieve System 2 reasoning, it is essential to establish a foundation for logical deduction and abstract thinking by capturing human intentions. Humans inherently perform System 2 reasoning through logical and abstract processes, and human intentions contain these crucial elements. Therefore, learning from human intentions, rather than merely imitating human actions, can provide a strong basis for advancing System 2 reasoning. Benchmarks like ARC [14], mini-ARC [30], 1D-ARC [66], ConceptARC [42], LARC [2], and MC-LARC [53] enable the collection of human behavior data to train AI models. Platforms like ARCreate Playground [31], O2ARC [52], ARC-Interactive [56], and ARC-Game [12] are useful for collecting such data. Additionally, environments like ARCLE [34] provide settings where models can learn from the collected data to improve reasoning and adaptability.
@@
These approaches are particularly effective for learning intentions from human trajectories in ARC tasks, where understanding and adapting to human strategies are crucial for success. For instance, in a task where the objective is to align a grid diagonally, Topic Modeling could help AI identify repeated rotations and flips as strong indicators of diagonal alignment strategies used by humans. Sequential Pattern Mining can further enhance this understanding by detecting recurring action sequences, such as Rotate(CW) followed by Flip(Horizontal), that consistently lead to correct outcomes. By identifying these patterns, the AI can begin to predict successful strategies and apply them to other, similarly structured tasks. Hidden Markov Models take this reasoning process a step further by modeling the transitions between actions and hidden goals, allowing the AI to generalize learned strategies across a wider variety of novel tasks. Incorporating these methods enables AI models to develop a deeper understanding of human reasoning, enhancing their logical deduction and adaptability, which are fundamental components of achieving System 2 reasoning capabilities. Thus, the combination of these techniques with data collected through platforms like O2ARC [52] not only improves AI’s ability to perform abstract reasoning but also strengthens its potential for handling complex, multi-step tasks in dynamic environments.
@@ 4.2 Combining Symbolic and Neural Models Improves Abstraction and Generality
Achieving System 2 reasoning requires models to move beyond simple pattern recognition and incorporate logical, structured reasoning capabilities. One promising approach to achieving this is through hybrid models that combine neural networks with symbolic reasoning. Neural networks excel at learning patterns from data, while symbolic reasoning provides interpretable, rule-based frameworks for logical deduction. By integrating these two methodologies, it is possible to enhance the ability to reason abstractly and generalize to new, unseen tasks, which are essential for System 2 reasoning.
Several methods have been developed to combine symbolic and neural models, enhancing generality and reasoning abilities.
Neuro-Symbolic Programming (NSP) [45] integrates the pattern recognition strengths of neural networks with the explicit logic of symbolic programming. In this approach, the neural network learns patterns from data, while the symbolic component generates interpretable code that can be applied to reasoning tasks. This hybrid method enables the handling of complex tasks requiring both learning from examples and reasoning over abstract concepts, such as program synthesis and abstract reasoning.
@@
In ARC tasks, hybrid neuro-symbolic models are well-suited for generalizing from a small number of examples and applying symbolic rules to solve new tasks. For example, NSP can learn how humans apply transformations like rotations and flips to grids, and then generalize these transformations using symbolic logic. ILP could infer rules based on observed task patterns, while symbolic regression might identify underlying mathematical relationships between grid elements. By combining neural networks and symbolic reasoning, these hybrid models can reason more effectively about the structure of tasks, enhancing their abstract reasoning capabilities and enabling more generalize to various tasks, key aspects of System 2 reasoning.
@@ 4.3 Meta-Learning Facilitates Adaptation to Unfamiliar Environments
System 2 reasoning requires AI systems to adapt quickly to new and unfamiliar tasks, a process that can be supported by meta-learning. Unlike traditional learning methods that are focused on task-specific optimization, meta-learning trains models to "learn how to learn,"making them more adaptable to various environments. By facilitating faster adaptation and improved generalization, meta-learning methods enable AI systems to respond effectively to novel challenges, which is crucial for achieving System 2 reasoning in tasks that require flexible thinking and logical deduction.
@@
Prototypical Networks [54] are designed for few-shot learning, where models classify new examples by comparing them to prototype representations. By learning these prototypes, AI systems can generalize across tasks with minimal data, making them efficient in environments that require structural reasoning. Prototypical networks are especially beneficial for System 2 reasoning, where recognizing and adapting to new patterns is key.
@@
In the context of ARC tasks, where multiple transformations may be required to achieve a goal, meta-learning can significantly enhance adaptability. For instance, in a task where the objective is to align objects diagonally, RNP could track the sequence of transformations and adjust reasoning based on temporal patterns. Prototypical networks could help the system recognize structural patterns involved in diagonal alignment, facilitating generalization across tasks. Meta-RL would provide the framework for balancing long-term planning with real-time reasoning, enabling the system to dynamically adjust its strategy based on feedback. By leveraging meta-learning, AI models can enhance their reasoning and adaptation capabilities, bringing them closer to the level of System 2 reasoning.